A cooling fan is turned off when it is running at 850 rev/min. It turns 1250 revolutions before it comes to a stop. (a) What was the fan’s angular acceleration, assumed constant? (b) How long did it take the fan to come to a complete stop?
Answer #1

Initial rotational speed of the fan = N1 = 850 rev/min

Initial angular speed of the fan = \omega 1

W = 27 Ni 60

2 (850) = الند 60

\omega1 = 89.01 rad/s

Final angular speed of the fan = \omega 2 = 0 rad/s (Comes to a stop)

Number of revolutions the fan makes before coming to a stop = n = 1250 rev

Angle through which the fan rotates before coming to a stop = \theta

\theta = 2\pin

\theta = 2\pi(1250)

\theta = 7853.98 rad

Angular acceleration of the fan = \alpha

\omega22 = \omega 12 + 2\alpha\theta

(0)2 = (89.01)2 + 2\alpha(7853.98)

\alpha = -0.5044 rad/s2

Negative as it is deceleration.

Time taken for the fan to come to a complete stop = T

\omega2 = \omega 1 + \alpha T

0 = 89.01 + (-0.5044)T

T = 176.5 sec

a) Angular acceleration of the fan = -0.5044 rad/s2

b) Time taken for the fan to come to a complete stop = 176.5 sec